OpenStudy (anonymous):

4 micrograms of a drug is administered to a patient. The amount of drug left in the blood after a certain period if time is given by the formula: A=4*(0.87)^t where t is the time(hours) and A is the amount of drug left. Algebraically determine how many hours after the drug has been administered that the amount of drug in the blood in the blood is halved.

OpenStudy (anonymous):

you'll have to solve this equation: 2 = 4*(0.87)^t

OpenStudy (anonymous):

1/2=(0.87)^t ln(1/2) = t*ln(0.87)

OpenStudy (anonymous):

sorry, but what is the "In" in the last equation?